National Advertising just paid a dividend of D0 = $0.75 per share, and that dividend is

expected to grow at a constant rate of 6.50% per year in the future. The company's

beta is 1.25, the required return on the market is 10.50%, and the risk-free rate is

4.50%. What is the company's current stock price?

Orwell Building Supplies' last dividend was $1.75. Its dividend growth rate is

expected to be constant at 25% for 2 years, after which dividends are expected to grow

at a rate of 6% forever. Its required return (rs) is 12%. What is the best estimate of the

current stock price?

*****Please show work on EXCEL spreadsheet, please show EXCEL formulas****
Answer #1

Cost of Equity using CAPM method =Risk free rate +Beta*(Market Return -Risk free rate) =4.5%+1.25*(10.5%-4.50%) =12%
Current Stock Price =D0*(1+growth)/(Cost of Equity-growth) =0.75*(1+6.5%)/(12%-6.5%) =14.52

Using Excel formula

D0 1.75
Growth 1 for 2 years 25%
Growth 2 after 2 years 6%
A B C D
Year 0 1 2 3
1 Excel formula=1.75*1.25 Excel formula=1.75*1.25^2 Excel formula=1.75*1.25^2*1.06
2 Dividend 1.75 2.1875 2.734375 2.8984375
3 Terminal Value 48.30729167 (Formula=1.75*1.25^2*1.06/(12%-6%)
4 Total 2.1875 51.04166667
Price if Stock $42.64 Excel formula=NPV(12%,B4:C4)